OpenStudy (anonymous):

Solve the formula for the given letter. Assume all variables represent nonnegative numbers. E=kc^2, for c. it's e/k, I just cant figure out whether it's positive or negative and with or without a sqrt

OpenStudy (anonymous):

Assuming you have: \[E=kc^2\] Divide each side by k: \[\frac{E}{k}=\frac{kc^2}{k}\] \[\frac{E}{k}=c^2\] Take the square Root on each side: \[c=\sqrt{\frac{E}{k}}\]